A stone has a mass of 5kg. If it takes the stone 4 sec to fall freely to the ground, calculate it's initial height above the ground, and the force with which the stone hit the ground.


1
Expert's answer
2021-12-15T11:18:37-0500

"\begin{array}{l}
m=5 kg \\
t=4 sec \\
h=\frac{g t^{2}}{2}=\frac{9,81 \cdot 4^{2}}{2}=78,48 m \\
F=m \cdot a \\
a=g \cdot t=9,81 \cdot 4=36,72 m / s ^{2} \\
F=36,72 \cdot 5=183,6 N
\end{array}"